OpenStudy (anonymous):

Bill performed an experiment in which he spun a spinner divided into four equal parts. Each part had a different color. The results are shown below. Color Times Spun red 19 blue 35 green 38 yellow 22 Based on the experiment above, predict the number of times the arrow will stop on the green part of the spinner if he performs the experiment a total of 342 times. A. 63 B. 72 C. 114 D. 75

OpenStudy (anonymous):

\[\frac{38}{19+35+38+22}=\frac{1}{3} \]\[\frac{342}{3}=114 \]

OpenStudy (anonymous):

Thank you for the medal.